# Plug-and-play imaging with model uncertainty quantification in radio astronomy

### 12 Dec 2023 v1

@@ Abstract
Plug-and-Play (PnP) algorithms are appealing alternatives to proximal algorithms when solving inverse imaging problems. By learning a Deep Neural Network (DNN) behaving as a proximal operator, one waives the computational complexity of optimisation algorithms induced by sophisticated image priors, and the sub-optimality of handcrafted priors compared to DNNs. At the same time, these methods inherit from the versatility of optimisation algorithms allowing to minimise a large class of objective functions. Such features are highly desirable in radio-interferometric (RI) imaging in astronomy, where the data size, the ill-posedness of the problem and the dynamic range of the target reconstruction are critical. In a previous work, we introduced a class of convergent PnP algorithms, dubbed AIRI, relying on a forward-backward backbone, with a differentiable data-fidelity term and dynamic range-specific denoisers trained on highly pre-processed unrelated optical astronomy images. Here, we show that AIRI algorithms can benefit from a constrained data fidelity term at the mere cost of transferring to a primal-dual forward-backward algorithmic backbone. Moreover, we show that AIRI algorithms are robust to strong variations in the nature of the training dataset: denoisers trained on MRI images yield similar reconstructions to those trained on astronomical data. We additionally quantify the model uncertainty introduced by the randomness in the training process and suggest that AIRI algorithms are robust to model uncertainty. Eventually, we propose an exhaustive comparison with methods from the radio-astronomical imaging literature and show the superiority of the proposed method over the state-of-the-art.
@@ I Introduction
Plug-and-play (PnP) approaches [9, 10, 11, 12] propose to circumvent the generalization difficulties of end-to-end methods by learning the proximal operator of an implicit prior unrelated to the measurement operator $H$ in (1). The resulting learned operator is then plugged in an optimisation algorithm, allowing to fully split the reconstruction task, alternating between steps enforcing data-fidelity (1), and steps enforcing an implicit prior. As a result, PnP algorithms are immune to variations in the sensing procedure at test time [13, 14], thus facilitating straightforward real-world applications. The performance of the resulting algorithm is closely tied to the denoiser’s quality, which accounts for the widespread adoption of denoising DNNs in PnP algorithms. However, despite DNNs’ proficiency in denoising tasks, their integration into PnP algorithms can introduce significant uncertainty into the reconstruction process. This underscores the need to quantify model-related uncertainty in images reconstructed using PnP algorithms.
@@
Our results show that adopting a constrained approach improves reconstructions over the current state-of-the-art, and that the proposed PnP algorithm is robust to the nature of images present in the training dataset. In particular, we show that our denoisers can be trained on MRI images and still perform well at test time on radio images, in simulated and real data settings. Eventually, we show that the training several denoisers allows to define a meaningful metric for quantifying model-based uncertainty in the reconstructions.
@@ II-A From splitting to PnP algorithms
When adopting a variational approach, the ill-posed problem (1) is relaxed in a well-posed problem where the sought image solution is assumed to satisfy both a good fit to the measurements, as well as some prior image model (such as smoothness or sparsity). Concretely, these relaxations take the form
@@
In the case where $f$ is differentiable, i.e. $\partial f=\{\nabla f\}$, a celebrated algorithm for solving (2) is the forward-backward (FB) splitting algorithm, writing as
@@
In the case where $f$ is not differentiable, i.e. $\partial f\neq\{\nabla f\}$, one needs to resort to algorithms slightly more complex than (3). In this setting, an algorithm fully leveraging the structure of (2) is the primal-dual (PD) algorithm [24], writing as
@@
It thus appears that the prior model, which is encoded in $A$ in (2), only percolate through the resulting algorithms at the level of the operator $\operatorname{D}$.
@@
In their seminal paper introducing Plug-and-Play (PnP) algorithms [9], the authors noticed that replacing the resolvent operator by a Gaussian denoiser allowed to improve the results over traditional handcrafted priors. This has since then been confirmed by a number of recent works leveraging the power of deep neural networks (DNNs) for denoising [10, 11, 13].
@@ II-C Related works
We eventually mention that similar approaches have been proposed in the literature. In [26, 27], the authors adopt a similar regularisation to ours to ensure convergence of the PnP algorithm, but they further rely on a differentiable DNN that allows to explicitly compute an associated loss function. Other approaches have been proposed in the literature for controling the stability of PnP algorithms [13, 14, 28, 26, 27, 29, 30] but they fall beyond the scope of the current article. While the original work and a large part of the PnP literature [9, 11] focused on the ADMM splitting algorithm [31], such procedure can be applied to any proximal splitting algorithm, and in particular, as will be of interest in this work, to the PD algorithm. We stress that the PnP variant of the PD algorithm has been used in a number of works [32, 10, 33, 34].
@@ III-A Constrained and unconstrained approaches
Such choice for $f$ corresponds to the most common one in PnP algorithms. In this case, since $f$ is differentiable, the PnP-FB algorithm can be used.
Motivated by compressed-sensing theory, another popular strategy for solving (1) when the sought solution is sparse (i.e. in radio astronomy) suggests to use a constrained data-fidelity term. In this case,
@@
where $\iota_{\mathcal{B}(z,\varepsilon)}(\cdot)$ denotes the indicator [^1] The indicator $\iota_{C}$ of a closed convex set $C$ is the function defined as $\iota_{C}(x)=0$ if $x\in C$ and $\iota_{C}(x)=+\infty$ otherwise. function of $\mathcal{B}(z,\varepsilon)$, the $\ell_{2}$ ball of radius $\varepsilon$ and centered in $z$. Since $f$ is not differentiable, the FB algorithm (3) cannot be applied, but the PD algorithm (4) offers an alternative.
While the two choices for $f$ lead to different algorithms and different problems (2), they lead to the same solution for the same $A$ if the hyperparameters of the problem are chosen appropriately. However, setting the hyperparameter $\varepsilon$ in (8) is often easier than tuning the regularisation of an unconstrained formulation. In the case of Gaussian noise for $n$ in (1) for instance, $\varepsilon$ corresponds to the variance of a chi-squared distribution, that can be computed from the (rather easily estimated) standard deviation of the Gaussian noise in the measurements. Furthermore, algorithms relying on the constrained formulation have shown to outperform their unconstrained counterparts. Algorithm (3) with AIRI denoisers plugged in is abbreviated as unconstrained AIRI (uAIRI). In contrast, algorithm (4) with AIRI denoisers plugged in is abbreviated as constrained AIRI (cAIRI).
@@ III-B Generating synthetic high dynamic range images
A distinctive characteristic of radio images is their high dynamic range. We define the dynamic range of an image as the ratio between its brightest and faintest non-zero pixels. While for 8-bit encoded natural images this ratio is of the order of $1/(1/255)=255$, for radio images, this value is typically $\gtrsim 10^{4}$. We stress that this is not a result of the encoding strategy of the measurements, but a characteristic of radio images, where some important features in radio images tend to have very faint intensity (for instance radio plumes or radio relics, see [35, 36]). As a consequence, radio images are usually displayed in logarithmic scale. In the remainder of this work we will use the following explicit transform $\operatorname{rlog}_{a}(x)=x_{\text{max}}\log_{a}(ax/x_{\text{max}}+1)$, which maps an image $x$ with dynamic range $a\gg 1$ and pixel values in $[x_{\text{max}}/a;x_{\text{max}}]$ to a low-dynamic range image $\operatorname{rlog}_{a}(x)$, with dynamic range $\log_{a}(2)$ and pixel values in the range $[x_{\text{max}}\log_{a}(2);x_{\text{max}}]$.
In the absence of a clean RI image datasets, we proposed in [22] a dataset of simulated high-dynamic range images and showed that such dataset could serve as a basis for training PnP denoisers, as well as end-to-end DNNs to solve the inverse problem (1). To simulate high dynamic range images, we applied the inverse of $\operatorname{rlog}_{a}$ to the images of interest. Denoting this inverse $\operatorname{rexp}_{a}$ defined as
@@
In order to investigate the influence of the nature of the images used in the training of the denoiser, we simulate a dataset of MRI images with similar dynamic range to that of RI images. The second dataset we study in this article contains MRI images extracted from the fastMRI single-coil knee images [6], where patches of each image are combined together in order to break symmetries and to increase image size. This exponentiation strategy allows us to generate 5000 MRI images of size 512$\times$512 with high dynamic range. Both the astronomical images and the MRI images were preprocessed with a blind image restoration network (SCUNet [37]) to remove noise and artefacts prior to exponentiation.
@@ III-C Dynamic range estimation and network selection
The performance of PnP algorithms strongly depends on the choice of the training noise level $\sigma$ in (6) [10, 14, 38]. Since the maximum pixel intensities of the images in the training dataset have been normalised to $1$, the value $\sigma$ defines the inverse of the dynamic range for training. A heuristic rule to choose this value has been established [22], which consists in setting it equal to the inverse of signal-to-noise ratio of the observed data, itself informing the target dynamic range of the reconstruction. The inverse of signal-to-noise ratio of the observed data can be estimated from the properties of $H$ and $e$ in (1) as:
@@
We thus propose to solve this equation numerically for the exponentiation parameter $a$ at different training noise levels we aim at using in our experiments. The $\sigma_{0}$ is selected such that the intensities of $1\%$ non-zero pixels in corresponding dataset are below this value.
@@ III-D Denoiser realisation and model uncertainty quantification
Similarly to [45], we instead propose in this work to investigate the epistemic uncertainty in the reconstruction, induced by variations in the denoiser $\operatorname{D}$ (and thus implicitly in the regularisation model characterised by $A$). In fact, while the choice of $A$ in (2) is cornerstone in the reconstruction quality, its relationship with the denoiser $\operatorname{D}$ in the case of PnP algorithms remains unclear since no closed form for $A$ is available in general, not to mention the high number of parameters and the (by nature) imperfect training procedure of $\operatorname{D}$.
@@
Similar approaches to ours were proposed in [47, 45]. Albeit minor differences, the unfolded algorithm from [45] also yields a set of estimates $\operatorname{Sol}$; the authors however train a single model the weights of which can be sampled, while we train a set of denoisers in a traditional fashion, and in their case, no link to a variational inclusion problem (2) can be made. We note that a large amount of works have studied aleatoric and epistemic uncertatinties in the context of end-to-end supervised learning with Bayesian neural networks [48, 49, 50, 51, 52, 53].
@@ IV-A Test dataset and evaluation metrics
Our test dataset consists in 3 radio images (3c353, Hercules A, and Centaurus A, see [22] for more details) of size $N=$ 512$\times$512. Each image has peak value normalised to 1, and a dynamic range slightly above $10^{4}$. For each image, we generate 4$\times$5 RI measurement vectors from different operators $H$ as (1). Each operator $H$ corresponds to a simulated MeerKAT telescope [54] acquisition simulation with specific pointing direction and observation time. We consider five distinct pointing directions chosen at random, and for each direction, we vary the observation time with $\Delta T$ selected from the set $\{1\text{h},2\text{h},4\text{h},8\text{h}\}$. This leads to a total of 4$\times$5 inverse problems for each groundtruth image, i.e. a total of 4$\times$15 inverse problems with $z\in\mathbb{C}^{M}$ containing $M\in\{4\times 10^{5},8\times 10^{5},1.6\times 10^{6},3.2\times 10^{6}\}$ data points, each value of $M$ being related to a value of $\Delta T$. Figure 1 (b) shows one of the test sampling patterns for a given pointing direction and observation time $\Delta T=4\text{h}$. The dynamic ranges of all the measurements fall in the range $[2.5\times 10^{3},1.4\times 10^{4}]$, and the average dynamic range of the measurements with $\Delta T=1,2,4,8$ are $2.9\times 10^{3}$, $3.9\times 10^{3}$, $5.6\times 10^{3}$ and $1.1\times 10^{4}$ respectively.
@@ IV-B Benchmark algorithms
We compare the performance of the proposed algorithms with several other methods: CLEAN [^2] https://gitlab.com/aroffringa/wsclean/ [20], a matching-pursuit algorithm which is the reference algorithm in the RI imaging community and is thus considered a meaningful baseline; a UNet [11] that was trained on the proposed synthetic datasets to solve the RI imaging problem in a supervised end-to-end fashion for each $\Delta T$; and the Sparsity Averaging Reweighted Analysis approaches uSARA [22] and SARA [^3] https://github.com/basp-group/Faceted-HyperSARA [17], which are, respectively, slightly evolved versions of the traditional FB (3) and PD algorithms (4) respectively, leveraging a redundant wavelet dictionary to impose simultaneous sparsity in multiple wavelet bases and a reweighting procedure to account for a nonconvex log-sum prior. uSARA and SARA have been exhaustively validated for RI imaging, where they have defined new state-of-the-art.
We further test our algorithms with the BM3D denoiser [55] in (3) and (4) instead of the proposed trained denoiser. We stress that BM3D is a non-trained denoiser, which allows to circumvent the lack of training data in RI imaging; however, it does not yield convergent PnP algorithms and suffer from slow inference time. The algorithm with a BM3D denoiser plugged into (3) is named uPnPBM3D and the one with BM3D plugged into (4) is named cPnPBM3D.

### 14 Dec 2023 v2

@@ Abstract
Plug-and-Play (PnP) algorithms are appealing alternatives to proximal algorithms when solving inverse imaging problems. By learning a Deep Neural Network (DNN) behaving as a proximal operator, one waives the computational complexity of optimisation algorithms induced by sophisticated image priors, and the sub-optimality of handcrafted priors compared to DNNs. At the same time, these methods inherit the versatility of optimisation algorithms allowing the minimisation of a large class of objective functions. Such features are highly desirable in radio-interferometric (RI) imaging in astronomy, where the data size, the ill-posedness of the problem and the dynamic range of the target reconstruction are critical. In a previous work, we introduced a class of convergent PnP algorithms, dubbed AIRI, relying on a forward-backward algorithm, with a differentiable data-fidelity term and dynamic range-specific denoisers trained on highly pre-processed unrelated optical astronomy images. Here, we show that AIRI algorithms can benefit from a constrained data fidelity term at the mere cost of transferring to a primal-dual forward-backward algorithmic backbone. Moreover, we show that AIRI algorithms are robust to strong variations in the nature of the training dataset: denoisers trained on MRI images yield similar reconstructions to those trained on astronomical data. We additionally quantify the model uncertainty introduced by the randomness in the training process and suggest that AIRI algorithms are robust to model uncertainty. Finally, we propose an exhaustive comparison with methods from the radio-astronomical imaging literature and show the superiority of the proposed method over the current state-of-the-art.
@@ I Introduction
Plug-and-play (PnP) approaches [9, 10, 11, 12] propose to circumvent the generalization difficulties of end-to-end methods by learning the proximal operator of an implicit prior unrelated to the measurement operator $H$ in (1). The resulting learned operator is then plugged into an optimisation algorithm, allowing the reconstruction task to be split, alternating between steps enforcing data-fidelity (1), and steps enforcing an implicit prior. As a result, PnP algorithms are immune to variations in the sensing procedure at test time [13, 14], thus facilitating straightforward real-world applications. The performance of the resulting algorithm is closely tied to the denoiser’s quality, which accounts for the widespread adoption of denoising DNNs in PnP algorithms. However, despite DNNs’ proficiency in denoising tasks, their integration into PnP algorithms can introduce significant uncertainty into the reconstruction process. This underscores the need to quantify model-related uncertainty in images reconstructed using PnP algorithms.
@@
Our results show that adopting a constrained approach improves reconstructions over the current state-of-the-art, and that the proposed PnP algorithm is robust to the nature of images present in the training dataset. In particular, we show that our denoisers can be trained on MRI images and still perform well at test time on radio images, in simulated and real data settings. Finally, we show that training several denoisers allows the definition of a meaningful metric for quantifying model-based uncertainty in the reconstructions.
@@ II-A From splitting to PnP algorithms
When adopting a variational approach, the ill-posed problem (1) is relaxed to a well-posed problem where the sought image solution is assumed to satisfy both a good fit to the measurements, as well as some prior image model (such as smoothness or sparsity). Concretely, these relaxations take the form
@@
In the case where $f$ is differentiable, i.e. $\partial f=\{\nabla f\}$, a celebrated algorithm for solving (2) is the forward-backward (FB) splitting algorithm, written as
@@
In the case where $f$ is not differentiable, i.e. $\partial f\neq\{\nabla f\}$, one needs to resort to slightly more complex algorithms than (3). In this setting, an algorithm fully leveraging the structure of (2) is the primal-dual (PD) algorithm [24], written as
@@
It thus appears that the prior model, which is encoded in $A$ in (2), only percolates through the resulting algorithms at the level of the operator $\operatorname{D}$.
@@
In their seminal paper introducing Plug-and-Play (PnP) algorithms [9], the authors noticed that replacing the resolvent operator by a Gaussian denoiser enabled improvements of the results compared to traditional handcrafted priors. This has since then been confirmed by a number of recent works leveraging the power of deep neural networks (DNNs) for denoising [10, 11, 13].
@@ II-C Related works
We note that similar approaches have been proposed in the literature. In [26, 27], the authors adopt a similar regularisation to the one we are proposing to ensure convergence of the PnP algorithm, but they further rely on a differentiable DNN that allows explicit computation of an associated loss function. Other approaches have been proposed in the literature for controlling the stability of PnP algorithms [13, 14, 28, 26, 27, 29, 30] but they fall beyond the scope of the current article. While the original work and a large part of the PnP literature [9, 11] focused on the ADMM splitting algorithm [31], such procedures can be applied to any proximal splitting algorithm, and in particular, as will be of interest in this work, to the PD algorithm. We stress that the PnP variant of the PD algorithm has been used in a number of works [32, 10, 33, 34].
@@ III-A Constrained and unconstrained approaches
Such a choice for $f$ corresponds to the most common one in PnP algorithms. In this case, since $f$ is differentiable, the PnP-FB algorithm can be used.
Motivated by compressed-sensing theory, another popular strategy for solving (1) when the sought solution is sparse (i.e. in radio astronomy) involves using a constrained data-fidelity term. In this case,
@@
where $\iota_{\mathcal{B}(z,\varepsilon)}(\cdot)$ denotes the indicator [^1] The indicator $\iota_{C}$ of a closed convex set $C$ is the function defined as $\iota_{C}(x)=0$ if $x\in C$ and $\iota_{C}(x)=+\infty$ otherwise. function of $\mathcal{B}(z,\varepsilon)$, the $\ell_{2}$ ball of radius $\varepsilon$ and centred in $z$. Since $f$ is not differentiable, the FB algorithm (3) cannot be applied, but the PD algorithm (4) offers an alternative.
While the two choices for $f$ lead to different algorithms and different problems (2), they lead to the same solution for the same $A$ if the hyperparameters of the problem are chosen appropriately. However, setting the hyperparameter $\varepsilon$ in (8) is often easier than tuning the regularisation of an unconstrained formulation. In the case of Gaussian noise for $n$ in (1) for instance, $\varepsilon$ corresponds to the variance of a chi-squared distribution, that can be computed from the (rather easily estimated) standard deviation of the Gaussian noise in the measurements. Furthermore, algorithms relying on the constrained formulation have been shown to outperform their unconstrained counterparts [22]. Algorithm (3) with AIRI denoisers plugged in is abbreviated as unconstrained AIRI (uAIRI). In contrast, algorithm (4) with AIRI denoisers plugged in is abbreviated as constrained AIRI (cAIRI).
@@ III-B Generating synthetic high dynamic range images
In the absence of clean RI image datasets, we proposed in [22] to generate a dataset of synthetic RI images to serve as a basis for training PnP denoisers as well as end-to-end DNNs to solve the inverse problem (1). A distinctive characteristic of radio images is their high dynamic range, where we define the dynamic range of an image as the ratio between its brightest and faintest non-zero pixels. While for 8-bit encoded natural images this ratio is of the order of $1/(1/255)=255$, for radio images, this value is typically $\gtrsim 10^{4}$. We stress that this is not a result of the encoding strategy of the measurements, but a characteristic of radio images, where some important features in radio images tend to have very faint intensity (for instance radio plumes or radio relics, see [35, 36]). As a consequence, radio images are usually displayed in logarithmic scale. In the remainder of this work we will use the following explicit transform $\operatorname{rlog}_{a}(x)=x_{\text{max}}\log_{a}(ax/x_{\text{max}}+1)$, which maps an image $x$ with dynamic range $a\gg 1$ and pixel values in $[x_{\text{max}}/a;x_{\text{max}}]$ to a low-dynamic range image $\operatorname{rlog}_{a}(x)$, with dynamic range $\log_{a}(2)$ and pixel values in the range $[x_{\text{max}}\log_{a}(2);x_{\text{max}}]$.
To simulate high dynamic range images, we apply the inverse of $\operatorname{rlog}_{a}$ to the images of interest. Denoting this inverse $\operatorname{rexp}_{a}$ defined as
@@
In order to investigate the influence of the nature of the images used in the training of the denoiser, we simulate a dataset of MRI images with similar dynamic range to that of RI images. The second dataset we study in this article contains MRI images extracted from the fastMRI single-coil knee images [6], where patches of each image are combined in order to break symmetries and to increase image size. This exponentiation strategy allows us to generate 5000 MRI images of size 512$\times$512 with high dynamic range. Both the astronomical images and the MRI images were preprocessed with a blind image restoration network (SCUNet [37]) to remove noise and artefacts prior to exponentiation.
@@ III-C Dynamic range estimation and network selection
The performance of PnP algorithms strongly depends on the choice of the training noise level $\sigma$ in (6) [10, 14, 38]. Since the maximum pixel intensities of the images in the training dataset have been normalised to $1$, the value $\sigma$ defines the inverse of the dynamic range for training. A heuristic rule to choose this value has been established [22], which consists of setting it equal to the inverse of signal-to-noise ratio of the observed data, itself informing the target dynamic range of the reconstruction. The inverse of signal-to-noise ratio of the observed data can be estimated from the properties of $H$ and $e$ in (1) as:
@@
We thus propose to solve this equation numerically for the exponentiation parameter $a$ at different training noise levels we aim at using in our experiments. The $\sigma_{0}$ is selected such that the intensities of $1\%$ non-zero pixels in the corresponding dataset are below this value.
@@ III-D Denoiser realisation and model uncertainty quantification
Similarly to [45], we instead propose in this work to investigate the epistemic uncertainty in the reconstruction, induced by variations in the denoiser $\operatorname{D}$ (and thus implicitly in the regularisation model characterised by $A$). In fact, while the choice of $A$ in (2) is a cornerstone of the reconstruction quality, its relationship with the denoiser $\operatorname{D}$ in the case of PnP algorithms remains unclear since no closed form for $A$ is available in general, not to mention the high number of parameters and the (by nature) imperfect training procedure of $\operatorname{D}$.
@@
Similar approaches to ours were proposed in [47, 45]. In particular, the unfolded algorithm from [45] also yields a set of estimates $\operatorname{Sol}$; the authors however train a single model the weights of which can be sampled, while we train a set of denoisers in a traditional fashion, and in their case, no link to a variational inclusion problem (2) can be made. We note that previous works have studied aleatoric and epistemic uncertainties in the context of end-to-end supervised learning with Bayesian neural networks [48, 49, 50, 51, 52, 53].
@@ IV-A Test dataset and evaluation metrics
Our test dataset consists of 3 radio images (3c353, Hercules A, and Centaurus A, see [22] for more details) of size $N=$ 512$\times$512. Each image has the peak value normalised to 1, and a dynamic range slightly above $10^{4}$. For each image, we generate 4$\times$5 RI measurement vectors from different operators $H$ as (1). Each operator $H$ corresponds to a simulated MeerKAT telescope [54] acquisition with specific pointing direction and observation time. We consider five distinct pointing directions chosen at random, and for each direction, we vary the observation time with $\Delta T$ selected from the set $\{1\text{h},2\text{h},4\text{h},8\text{h}\}$. This leads to a total of 4$\times$5 inverse problems for each groundtruth image, i.e. a total of 4$\times$15 inverse problems with $z\in\mathbb{C}^{M}$ containing $M\in\{4\times 10^{5},8\times 10^{5},1.6\times 10^{6},3.2\times 10^{6}\}$ data points, each value of $M$ being related to a value of $\Delta T$. Figure 1 (b) shows one of the test sampling patterns for a given pointing direction and observation time $\Delta T=4\text{h}$. The dynamic ranges of all the measurements fall in the range $[2.5\times 10^{3},1.4\times 10^{4}]$, and the average dynamic range of the measurements with $\Delta T=1,2,4,8$ are $2.9\times 10^{3}$, $3.9\times 10^{3}$, $5.6\times 10^{3}$ and $1.1\times 10^{4}$ respectively.
@@ IV-B Benchmark algorithms
We compare the performance of the proposed algorithms with several other methods: CLEAN [^2] https://gitlab.com/aroffringa/wsclean/ [20], a matching-pursuit algorithm which is the reference algorithm in the RI imaging community and is thus considered a meaningful baseline; a UNet [11] that was trained on the proposed synthetic datasets to solve the RI imaging problem in a supervised end-to-end fashion for each $\Delta T$; and the Sparsity Averaging Reweighted Analysis approaches uSARA [22] and SARA [^3] https://github.com/basp-group/Faceted-HyperSARA [17], which are, respectively, slightly evolved versions of the traditional FB (3) and PD algorithms (4), leveraging a redundant wavelet dictionary to impose simultaneous sparsity in multiple wavelet bases and a reweighting procedure to account for a nonconvex log-sum prior. uSARA and SARA have been exhaustively validated for RI imaging, where they have defined new state-of-the-art.
We further test our algorithms with the BM3D denoiser [55] in (3) and (4) instead of the proposed trained denoiser. We stress that BM3D is a non-trained denoiser, which allows circumventing the lack of training data in RI imaging; however, it does not yield convergent PnP algorithms and suffer from slow inference time. The algorithm with a BM3D denoiser plugged into (3) is named uPnPBM3D and the one with BM3D plugged into (4) is named cPnPBM3D.